OpenStudy (anonymous):

The area of this cylinder's base is square inches. The height of the cylinder is 6 inches. What is the volume of the cone that has the same base area and height? Show the steps of your solution plaese help

OpenStudy (jack1):

Area... =? not enough info yet dude

OpenStudy (jack1):

volume of cylinder = A of B x Ht Volume of cone = Aof B x Ht x 1/3 therefore area of cone = 1/3 area of cylinder

OpenStudy (jack1):

sorry, typo, volume of cone = 1/3 volume of cylinder

OpenStudy (jack1):

without an answer for either the total volume of one of the shapes, the radius of the cone or the area of the base, can't help you further dude, sorry

OpenStudy (anonymous):

thank any way